# An engine not a camera: Measuring performative power of online search

### v1

@@ 4 Powermeter: Experiment in the wild
We designed an online experiment to measure the performativity gap on two popular online search platforms operated by Google and Bing. The experiment is built around a Chrome browser extension that modifies the arrangement of search result pages and records user click statistics in a privacy-preserving fashion. Thereby, the extension allows us to estimate the performativity gap over an organic set of clicks and arrangements without having control over the platforms’ underlying algorithm.
@@ 4.1 Browser extension
Powermeter is a Chrome browser extension designed to implement different counterfactual arrangements on Google search and Bing search. Extensions can add functionalities to the web-browser and change how a website is displayed to the user. We emphasize that Powermeter takes advantage of these capabilities, to hide or reorder, but never modify or add any content on the search page.
@@
Once activated, the extension triggers the experiment whenever the user enters a search query on either Google search or Bing search. This can be identified by monitoring the url string of the current tab. Before search results are loaded the extension immediately hides the content of the website, inspects the html document, randomly assigns the user to one of the experimental groups and then implements the respective changes to the website before making the page visible. The implementation of the counterfactuals is done by identifying the relevant items to hide or swap by their html class names or ids. We also add custom tags and event listeners to the identified elements that we can fall back on at a later stage. The entire setup of the experiment usually takes around 40 milliseconds. Hiding the html body of the website with the first possible Chrome event is crucial to avoid glitches in case of bad internet connection and make sure the control arrangement is not revealed to the participant. To ensure internal validity of our experiment, we also have to ensure a participant is not reassigned to a new experimental group when reloading a page, navigating between tabs or repeatedly entering the same search query. This is done by storing a hash of user ID and search query together with the assigned experimental group in the browser cache.
@@
Every participant is assigned a unique random number that serves as anonymous user ID upon installation of the extension. This user ID is stored in the local browser cache and persists throughout the experiment. Every time a click event on an element on the search page is registered, the click data is aggregated into a json object and sent to a database server hosted locally at out institution via a post request using the encrypted https protocol. This concerns information about the index of the clicked search result, the click element type, the page index, and the experimental group. In addition, statistics about the website such as the number of search results, the presence of ads and boxes, the number of candidate results, and the position of specialized search results are extracted from the website are recorded. The database server is built using the Microsoft .Net core framework and deployed within a docker container. The database access is rate limited and the Get endpoint of the database is key protected. We use a SQlite database that is mapped to persistent memory.
@@
The information that is stored with every click does not contain any personally identifiable information. While we record statistics and the position of the clicked element on the search page, we never store search queries or any information about the websites visited by the user. This is an intentional choice to preserve user privacy, and demonstrate that valuable insights can be gathered without privacy invasive data collection. The experiment went through an internal approval procedure and the privacy policy can be found on our website. [^4] The project website can be found at: https://powermeter.is.tue.mpg.de/ We also provide a copy in the appendix.
4.2 Experimental groups
We implement six different counterfactual arrangements, summarized in Table 1. We refer to Figure 2 for the terminology used to refer to individual elements on the general search page. The first three arrangements $a_{1}$$a_{3}$ concern the reordering of organic search results, leaving the other elements on the website untouched. The arrangements $a_{4}$ and $a_{6}$ perform modifications not directly concerning organic search results; Arrangement $a_{6}$ hides a specific element, called the Shopping box, appearing either in the right side panel or on top of the search results page, Arrangement $a_{4}$ hides the box together with all the Ads. Finally, Arrangement $a_{5}$ combines the latter change with a change in search result order. For Bing we only implemented the counterfactual $a_{1}$ to ensure statistical power despite data scarcity. A practical reason not to implement larger modifications is also users’ sensitivity to the resulting deterioration of quality. The Bing experiment of the European Commissions’s investigation had to be discontinued after one week for that exact reason. [^5] See Case T-612/17, supra 2, para 399 Furthermore, our minimal modifications are unlikely to have an immediate effect on user’s browsing behavior, or even be perceived by the user, which is important towards ensuring Assumption 1.
@@ 4.3 Onboarding
Participants were provided the link to the project website as an entry point. The website contains information about the experiment, the purpose of the study, an onboarding video, as well as the privacy policy of the extension. The extension itself is distributed through the official Chrome webstore and there is a button directly navigating the user to the store. We did not list the extension publicly to ensure participants are informed about the purpose of the study, and protect the integrity of our data. The installation follows the standard procedure of adding a browser extension to Chrome. The user has to give consent to access Google and Bing websites, as well as to use the storage API. The extension remains active until participants remove it from their web-browser, or until the experiment is stopped. The study participants are trusted individuals of different age groups and backgrounds, recruited by reaching out personally or via email. For each of the participants we observe all the search queries they perform while the extension is activated. We provide demographic statistics over our pool of participants in Figure 9 in the appendix.
@@
For each participant we ignored the clicks collected during the first four days after onboarding in order to avoid confounding by changes in behavior due to participation awareness (Keusch et al., 2022). We also removed clicks where the search elements could not be identified reliably for implementing the RCT to avoid selection bias towards the control group.
@@ 5 Empirical results
Using our Powermeter browser extension we collected click data from $85$ participants over the course of 4 months, from September 2023 until January 2024. This resulted in $56,971$ click events, and a total of $45,625$ clicks after preprocessing. Out of the clicks $98.9\%$ were registered on Google, and $1.1\%$ on Bing. Figure 8 in the appendix visualizes some aggregate statistics over the clicks collected. We will consider several subsets of these events for which we measure the performativity gap and algorithmic distortion. In the following we discuss the main insights from the collected data. For all plots, we provide bootstrap confidence bounds over 200 resamples.
@@ 5.1 Reordering search results
We first inspect the three counterfactual arrangements $a_{1}$, $a_{2}$, $a_{3}$. Recall that $c_{i}$ indexes search result $i$ in the order in which it appears under $a_{0}$. In Figure 3 we visualize the event probabilities $C=c_{i}$ for the control group (blue bars) and compare it to the respective probabilities under the three counterfactuals (orange bars). The three figures on the left show the results across Google search queries. Here the counterfactuals correspond to swapping the position of the first two results (left), the first and the third (middle) and the second and the third (right). The right figure shows the results evaluated on Bing search queries when the first two results are swapped. The lower figure visualizes the corresponding performativity gap $\delta^{i}(a)$ for $i=1,2,...,6$.
@@
We observe a consistently large effect of arrangement on clicks. Being down ranked by one position on Google search decreases average click through rate of $c_{1}$ from $43\%$ to $24\%$, resulting in $\delta^{1}(a_{1})=-0.19$ and an algorithmic distortion of $\beta=0.44$. Being down-ranked by 2 positions results in $\delta^{1}(a_{2})=-0.27$ and an average loss of more than $50\%$ of traffic. Note that a similar effect size has been reported in the case decision for the UK market for a two position shift (European Commission, 2017, 460), indicating that the estimate is robust across different populations. Finally, by down-ranking the second content element by one position we still observe a significant traffic reduction, corresponding to an algorithmic distortion of $\beta=0.39$.
@@
An interesting observation is also that for every counterfactual arrangement, the element shown first ends up getting most clicks on average, despite implementing a different ranking. Implying that all the rankings are performatively stable (Perdomo et al., 2020) with respect to the non-personalized reranking strategies considered. However, there are several indications of Google’s ranking $a_{0}$ reflecting relevance of search results better then the other arrangements. Namely, $c_{1}$ gets more clicks when displayed first, compared to other results in the same position, e.g., the first result under $a_{2}$ or $a_{3}$. A second indicator is that $c_{2}$ benefits from arrangement $a_{3}$; under the hypothesis that users consider search results in order this indicates that content item $c_{3}$ absorbs less clicks than $c_{1}$.
@@
Figure 3: Click through rate and performativity gap under the counterfactual arrangements $a_{1}$, $a_{2}$, $a_{3}$ for Google and the counterfactual arrangement $a_{1}$ for Bing, compared to the control arrangement $a_{0}$.
@@
For Bing the position effect seems to be even more pronounced, although confidence intervals are significantly larger in this case. We conjecture that this could be due to the average number of specialized search results and Ads present on the search page being larger on Bing, as this results in a larger spacial separation of search results. Statistics about elements present on the the search pages shown to our study participants are reported in Figure 8 in the Appendix.
@@ 5.2 Indirect effect of visually appealing elements
Next, consider the counterfactuals $a_{4}$ and $a_{6}$ that leave generic search results untouched and hide certain elements on the website. We first inspect the number of clicks these elements absorb if added to the page. We focus on Google search. In Figure 4 we compare the fraction of clicks going to generic search results, Ads, and Boxes for $a_{0}$, $a_{4}$, $a_{6}$. We plot the statistics across the aggregate queries (left panel), the subset of queries where Ads are present on the page under $a_{0}$ (middle panel) and the subset of queries where the box is present on the page under $a_{0}$ (right panel). We find the addition of boxes absorbs $22.4\%$ of the clicks on average across queries where it is present under $a_{0}$ and these clicks are mostly taken away form the generic search results. Similarly, Ads absorb close to $30\%$ of the clicks on average for queries where they are present. However, considering the overall number of clicks the effect is smaller since a large fraction of queries contains neither Ads nor Boxes.
@@
Figure 5: Effect of hiding box and swapping elements on the click probability of generic search results. Statistics are evaluated for the subset of queries for which box is naturally present on Google. The hashed bar shows the click probability under $a_{5}$ when top content is hidden and the first two elements are swapped. The right table reports the empirical measure of algorithmic distortion for different conducts, extracted form the results in the left figure.
@@
We now consider the combined conduct of adding the box and down-ranking an element. We constrain our focus on queries where the Shopping box is present under $a_{0}$, either in the center column or in the right sidebar. These are $3.2\%$ of all the events.
We show the corresponding click probabilities for the three first search results in Figure 5. In both figures the blue bars correspond to the control group $a_{0}$, and the red bars correspond to $a_{1}$. For these groups boxes are present on the page. In the left figure we investigate the effect of hiding boxes only and the orange bars correspond to arrangement $a_{4}$. In the right figure we investigate the effect of also hiding Ads, here the orange bars correspond to $a_{6}$. We find that when adding Boxes, all three content items loose a significant fraction of clicks, whereas Ads mostly take away from $c_{1}$. In the right figure we additionally show $a_{5}$ using the hatched bars (i.e., down-ranking the first element by one position if box is hidden). For $c_{1}$ the combined effect of adding Ads and Boxes on the click through rate is almost as large as the effect of down ranking the same item by one position. What we can see consistently is that combining the conduct of adding visually appealing elements on top of the page, and down-ranking a content item, has a combined effect that is larger than the effect of the two individual modifications alone. For element $c_{1}$ the measured distortion is reported in the table. The combined effect leads to a reduction of $25\%$ in clicks and an algorithmic distortion of $66\%$ when considering the effect of Ads/Boxes, and $53\%$ when only considering Boxes (comparing orange and red bar).
5.3 Factors that impact performativity
We perform additional investigations into what factors have a reinforcing effect on the performativity of the ranking. To this end, we inspect different subsets of Google queries and measure the performativity gap as well as algorithmic distortion for $c_{1}$ under the counterfactual $a_{1}$. First, we split the data across two different axes depending on whether Ads or boxes are present, and whether Specialized Search results (SSR) are present in between the first two results. The respective comparisons are visualized in the left and middle panel of Figure 6. We observe that the performativity gap in the presence of Ads and boxes is smaller, and about the same whether special search results are present in between search results. However, if we plot distortion we get a different picture, since the base click probability of content item $c_{1}$ across different splits can be very different for the three cases. We find that while Ads and Boxes have little effect on distortion, the presence of a specialized search result in between the swapped results tends to increase the effect of down-ranking on $c_{1}$.
@@ 5.3 Factors that impact performativity
For the second investigation we group the queries by the number of candidate search results available on Google. This number was extracted form the website where it appeared as a string on top of the page in the form ‘About 323’000’000 results (0.65 second)’. The right panel of Figure 6 shows algorithmic distortion for each percentile of the data. We see a clear trend that the performativity gap increases with the number of candidate results. We suspect this to be connected to the smaller relevance gap across results for queries with more potential results, leading to a higher influence of the arrangement on clicks. However, note that findings in this subsection are no causal claims.
@@ 6 Discussion
In this work we present a flexible experimental design, based on a browser extension, to investigate the effect of search algorithms on user clicks. The browser extension performs interventions at the level of display to emulate algorithmic updates, without access to the platform algorithm. We implement different counterfactuals, and measure their effect on clicks. In turn, we formally relate our quantitative findings to an instantiation of performative power measuring the platform’s ability to steer user clicks going out from its website.
@@
Next, we describe how our findings could fit into a broader anti-trust investigation. Take the Google Shopping case as an example. It is concerned with the ability of Google search to distort incoming traffic to a particular business’ website. Suppose, as a thought experiment, $80\%$ of the search referrals to the site come from Google Search. [^6] As part of the Google Shopping investigation (Section 7.2.4.1 in the decision, Table 24) it is reported that up to $82\%$ of incoming traffic to CS services in the European Economic Area (EEA) was mediated by Google search in 2012. Further, assume that $70\%$ of the referrals from Google happen while the service is ranked among the top two generic search results. Our estimates suggest that distortion of traffic at the first position can be as large as $66\%$ for small arrangement changes. Assuming for the second position the effect is $20\%$ smaller, giving a conservative average effect size of $\beta=0.8\cdot 0.66$. Multiplying the effect size by the fraction of incoming clicks it concerns, we get $0.8\cdot 0.7\cdot 0.9\cdot 0.66\approx 33\%$. This is the fraction of traffic to the site Google can influence.
@@
Turning this number into a conservative lower bound on performative power it offers an interpretable measure of power for an investigator to judge whether the use of this algorithmic lever should be a concern for competition and consumers or not. We can use the same logic to compare search engines, and assess the effectiveness of mandated remedies. One avenue for future work to strengthen such an argument is to rigorously argue about the external validity of empirical findings like ours.
@@
More broadly, we hope our work can serve as a blue print for how performative power can be used to guide digital market investigations, and how tools from computer science, causality, and performative prediction can inform ongoing legal debates related to the power of digital platforms. Our work is situated within a growing scholarship (c.f., Metaxa et al., 2021; Bandy, 2021; Dash et al., 2024) that takes advantage of the accessibility of digital markets for monitoring and regulating digital platforms. In addition to observational methods, we demonstrate how experiments can offer an additional powerful lever.
@@
From the perspective of computer science our work offers measurements and provides evidence of performativity in the context of online search, contributing empirical support to the study of performativity on online platforms. As its name suggests, a search engine acts as an engine steering consumption through its ranking algorithm, rather than a camera merely picturing candidate results—we borrow this analogy from MacKenzie (2008).

### v2

@@ 4 Powermeter: Experiment in the wild
We designed an online experiment to measure the performativity gap on two popular online search platforms operated by Google and Bing. The experiment is built around a Chrome browser extension that modifies the arrangement of search result pages and records user click statistics in a privacy-preserving fashion. The extension allows us to observe an organic set of user queries and click outcomes under different arrangements without having control over the platforms’ algorithm.
@@ 4.1 Browser extension
Browser extensions can add functionalities to the web-browser and change how a website is displayed to the user. Powermeter makes use of these functionalities to emulate algorithmic updates by implementing different counterfactual arrangements on Google search and Bing search. We emphasize that Powermeter only hides or reorders, but never modifies or adds any content on the search page.
@@
Once activated, the extension triggers the experiment whenever the user enters a search query on either Google search or Bing search. This can be identified by monitoring the url string of the current tab. Before search results are loaded the extension immediately hides the content of the website, inspects the html document, randomly assigns the user to one of the experimental groups and then implements the respective changes to the website before making the page visible. The implementation of the counterfactuals is done by identifying the relevant items to hide or swap by their html class names or ids. We also add custom tags and event listeners to the identified elements that we can fall back on at a later stage. The entire setup of the experiment usually takes around 40 milliseconds. This delay is far below what was found to be noticeable to users (Brutlag et al., 2008; Arapakis et al., 2014). Hiding the html body of the website with the first possible Chrome event is crucial to avoid glitches in case of bad internet connection and make sure the control arrangement is not revealed to the participant. To ensure internal validity of our experiment, we also have to ensure a participant is never reassigned to a new experimental group when reloading a page, navigating between tabs or repeatedly entering the same search query. This is done by storing a hash of user ID and search query together with the assigned experimental group in the browser cache.
@@
Every participant is assigned a unique random number that serves as anonymous user ID upon installation of the extension. This user ID persists throughout the experiment. Every time a click event on an element on the search page is registered, the click data is aggregated into a json object and sent to a database server hosted locally at out institution via a post request using the encrypted https protocol. This concerns information about the index of the clicked search result, the click element type, the page index, and the experimental group. In addition, statistics about the website such as the number of search results, the presence of ads and boxes, the number of candidate results, and the position of specialized search results are extracted from the website are recorded. The database server is built using the Microsoft .Net core framework and deployed within a docker container. The database access is rate limited and the Get endpoint of the database is key protected. We use a SQlite database that is mapped to persistent memory.
@@
The information that is stored with every click does not contain any personally identifiable information. While we record the position of the clicked element on the search page, we never store search queries or any information about the websites visited by the user. This is an intentional choice to preserve user privacy, and to demonstrate that valuable insights can be gathered without privacy invasive data collection. The experiment went through an internal approval procedure and the privacy policy can be found on our website. [^4] The project website can be found at: https://powermeter.is.tue.mpg.de/
@@
4.2 Experimental groups
We implement six different counterfactual arrangements, summarized in Table 1, each defining a treatment group. We refer to Figure 2 for the terminology used to refer to individual elements on the general search page. It equivalently applies to both, Google search and Bing. Arrangements $a_{1}-a_{6}$ are designed to emulate conservative variants of the conduct $a^{*}$ of interest, to inform a plausible lower bound on performative power. The first three arrangements $a_{1}$$a_{3}$ concern the reordering of organic search results, leaving the other elements on the website untouched. The arrangements $a_{4}$ and $a_{6}$ perform modifications not directly concerning organic search results: Arrangement $a_{6}$ hides a specific element, called the Shopping box, appearing either in the right side panel or on top of the search results page. Arrangement $a_{4}$ hides the box together with all the Ads. Finally, Arrangement $a_{5}$ combines the latter change with a change in search result order. For Bing we only implemented the counterfactual $a_{1}$ to ensure statistical power despite data scarcity. A practical reason not to implement larger modifications is also users’ sensitivity to the resulting deterioration of quality towards ensuring Assumption 1. The Bing experiment of the European Commission’s investigation had to be discontinued after one week for that exact reason. [^5] See Case T-612/17, supra 2, para 399 We made sure to avoid a similar failure point. Based on user feedback collected during an initial test round there was no indication that the modifications were even noticeable to users.
@@ 4.3 Onboarding
Participants were provided the link to the project website as an entry point. The website contains information about the experiment, the purpose of the study, an onboarding video, as well as the privacy policy of the extension. The extension itself is distributed through the official Chrome webstore and there is a button directly navigating the user to the item in the store. We did not list the extension publicly to ensure participants are informed about the purpose of the study, and protect the integrity of our data. The installation follows the standard procedure of adding a browser extension to Chrome. The user has to give consent to access Google and Bing websites, as well as to use the storage API. The extension remains active until participants remove it from their web-browser, or until the experiment is stopped. The study participants are trusted individuals of different age groups and backgrounds, recruited by reaching out personally or via email. We provide demographic statistics over our pool of participants in Figure 9 in the appendix.
@@
For each participant we ignored the clicks collected during the first four days after onboarding, as suggested by Keusch et al. (2022) in order to avoid potential confounding due to participation awareness. We also removed clicks where the search elements could not be identified reliably for implementing the RCT to avoid selection bias towards the control group.
@@ 5 Empirical results
Using our Powermeter browser extension we collected click data from $85$ participants over the course of 5 months, from September 2023 until January 2024. This resulted in $56,971$ click events, and a total of $45,625$ clicks after preprocessing. Out of the clicks $98.9\%$ were registered on Google, and $1.1\%$ on Bing. Figure 8 in the appendix visualizes some aggregate statistics over the clicks collected. We will consider several subsets of these events for which we measure the performativity gap and algorithmic distortion. In the following we discuss the main insights from the collected data. For all plots, we provide bootstrap confidence bounds over 200 resamples.
@@ 5.1 Reordering search results
We first inspect the three counterfactual arrangements $a_{1}$, $a_{2}$, $a_{3}$ concerning reranking. Recall that $c_{i}$ indexes search results in the order in which they appears under $a_{0}$. In Figure 3 we visualize the event probabilities $C=c_{i}$ for each search result $i=1,2,...,6$ under the control group (blue bars) and compare it to the respective probabilities under the three counterfactuals (orange bars). The figures on the left show the results across Google search queries. Here the counterfactuals correspond to swapping the position of the first two results (left), the first and the third (middle) and the second and the third (right). The right figure shows the results evaluated on Bing search queries when the first two results are swapped. The lower figure visualizes the corresponding performativity gap $\delta^{i}(a)$, corresponding to the change in clicks to item $i$ caused by the respective arrangement change.
@@
We observe a consistently large effect of arrangement on clicks. Being down ranked by one position on Google search decreases average click through rate of $c_{1}$ from $43\%$ to $24\%$, resulting in $\delta^{1}(a_{1})=-0.19$ and an algorithmic distortion of $\beta=0.44$. Being down-ranked by 2 positions results in $\delta^{1}(a_{2})=-0.27$ and an average loss of more than $50\%$ of traffic. Note that a similar effect size has been reported in the case decision for the UK market for a two position shift (European Commission, 2017, para 460), indicating that the estimate is robust across different user populations. Finally, by down-ranking the second content element by one position we still observe a significant traffic reduction, corresponding to an algorithmic distortion of $\beta=0.39$.
@@
An interesting observation is also that for every counterfactual arrangement, the element shown first ends up getting most clicks on average. Implying that all the rankings are close to performatively stable (Perdomo et al., 2020) with respect to the non-personalized reranking strategies considered. However, there are several indications of Google’s ranking $a_{0}$ reflecting relevance of search results better then the other arrangements. Namely, $c_{1}$ gets more clicks when displayed first, compared to other results displayed in the same position ($c_{2}$ corresponds to first result under $a_{2}$ or and $c_{3}$ to the first result under $a_{3}$). A second indicator is that $c_{2}$ benefits from arrangement $a_{3}$; under the hypothesis that users consider search results in order this indicates that content item $c_{3}$ absorbs less clicks than $c_{1}$.
@@
Figure 3: Click through rate and performativity gap for general search results $c_{1}$ to $c_{6}$ under the counterfactual arrangements $a_{1}$, $a_{2}$, $a_{3}$ for Google and the counterfactual arrangement $a_{1}$ for Bing, compared to the control arrangement $a_{0}$ (in blue).
@@
For Bing the position effect seems to be even more pronounced, although confidence intervals are significantly larger. We conjecture that this could be due to the average number of specialized search results and Ads present on the search page being larger on Bing. This results in a larger spacial separation of general search results and potentially larger display effects. Statistics about the type of elements present on the the search pages are reported in Figure 8 in the Appendix.
@@ 5.2 Indirect effect of visually appealing elements
Next, we consider the counterfactuals $a_{4}$ and $a_{6}$ that leave generic search results untouched and hide certain elements on the website. We first inspect the number of clicks these elements absorb if present on the page. We focus on Google search. In Figure 4 we compare the fraction of clicks going to generic search results, Ads, and Boxes for $a_{0}$, $a_{4}$, $a_{6}$. We plot the statistics across the aggregate queries (left panel), the subset of queries where Ads are present on the page under $a_{0}$ (middle panel) and the subset of queries where the box is present on the page under $a_{0}$ (right panel). We find the addition of boxes absorbs $22.4\%$ of the clicks on average across queries where it is present under $a_{0}$ and these clicks are mostly taken away form the generic search results. Similarly, Ads absorb close to $30\%$ of the clicks on average for queries where they are present. However, considering the overall number of clicks the effect is smaller since a large fraction of queries contains neither Ads nor Boxes.
@@
Figure 5: Effect of hiding box and swapping elements on the click probability of generic search results. Statistics are evaluated for the subset of queries for which box is naturally present. The hashed bar shows the click probability under $a_{5}$ when top content is hidden and the first two elements are swapped. The right table reports the empirical measure of algorithmic distortion for different conducts, extracted form the results in the left figure.
@@
We now consider the combined conduct of adding the box and down-ranking an element. We constrain our focus on queries where the Shopping box is present under $a_{0}$, either in the center column or in the right sidebar. These are $3.2\%$ of all the events. We show the corresponding click probabilities for the three first search results in Figure 5. In both figures the blue bars correspond to the control group $a_{0}$, and the red bars correspond to $a_{1}$. For these groups boxes are present on the page. In the left figure we investigate the effect of hiding boxes only and the orange bars correspond to arrangement $a_{4}$. In the right figure we investigate the effect of also hiding Ads, here the orange bars correspond to $a_{6}$. We find that when adding Boxes, all three content items loose a significant fraction of clicks, whereas Ads mostly take away from $c_{1}$. In the right figure we additionally show $a_{5}$ using the hatched bars (i.e., down-ranking the first element by one position if box is hidden). For $c_{1}$ the combined effect of adding Ads and Boxes on the click through rate is almost as large as the effect of down ranking the same item by one position. What we can see consistently is that combining the conduct of adding visually appealing elements on top of the page, and down-ranking a content item, has a combined effect that is larger than the effect of the two individual modifications alone. For element $c_{1}$ the measured distortion is reported in the table. The combined effect leads to a reduction of $25\%$ in clicks and an algorithmic distortion of $66\%$ when considering the effect of Ads/Boxes, and $53\%$ when only considering Boxes (comparing orange and red bar). We believe this to be the first time that quantitative insights into this combined conduct are made public.
5.3 Factors that impact performative power
We perform additional investigations into what factors have a reinforcing effect on the performativity of ranking. To this end, we inspect different subsets of Google queries and measure the performativity gap as well as algorithmic distortion for $c_{1}$ under the counterfactual $a_{1}$. First, we split the data across two different axes depending on whether Ads or boxes are present, and whether Specialized Search results (SSR) are present in between the first two generic search results. The respective comparisons are visualized in the left and middle panel of Figure 6. We observe that the performativity gap in the presence of Ads and boxes is smaller, and about the same whether special search results are present in between search results. However, if we plot distortion we get a different picture, since the base click probability of content item $c_{1}$ across different splits is different for the three cases. We find that while Ads and Boxes have little effect on distortion, the presence of a specialized search result in between the swapped results tends to increase the effect of down-ranking $c_{1}$.
@@ 5.3 Factors that impact performative power
For the second investigation we group the queries by the number of candidate search results available on Google. This number was extracted from the website where it appeared as a string on top of the page in the form ‘About 323’000’000 results (0.65 second)’. The right panel of Figure 6 shows algorithmic distortion for each percentile of the data. We see a clear trend that the performativity gap increases with the number of candidate results. We suspect this to be connected to the smaller relevance gap across results for queries with more potential results, leading to a higher influence of the arrangement on clicks. However, note that findings in this subsection are no causal claims.
@@ 6 Discussion
We presented a flexible experimental design, based on a browser extension, to investigate the effect of search algorithms on user clicks. The browser extension performs interventions at the level of display to emulate algorithmic updates, without access to the platform algorithm. We implement different counterfactuals relevant for the Google Shopping antitrust investigation, and provide quantitative insights into the causal effect on clicks. Theorem 1 formally relates these quantitative findings to an instantiation of performative power, measuring the platform’s ability to steer outgoing traffic from its general search service.
In a final step, we describe how our findings could fit with a broader anti-trust investigation potentially concerning effects spanning different markets. Take the Google Shopping case as an example. The case is concerned with the ability of Google search to distort incoming traffic to a business operating in the market of comparison shopping services by changing where it appears on Google search relative to its competitors. Establishing the causal link between arrangements on search and their effect on incoming traffic to a third party website composes into two steps: a) establish Google’s ability to steer outgoing traffic, b) quantify how much of the incoming traffic is mediated by Google search. The first step is a notion of power that experiments like ours operationalize, the second is a number that can readily be obtained from web traffic data. The overall performative power will be the product of the two factors. For putting this together, let’s work though the following thought experiment:
@@
Suppose, $80\%$ of the referrals to the competitor’s website come from Google Search. [^6] In 2022 up to $82\%$ of incoming traffic to Comparison Shopping Services in the European Economic Area was mediated by Google search, as reported in the Google Shopping case decision (Section 7.2.4.1, Table 24). Further, assume that $70\%$ of the referrals from Google happen while the website is ranked among the top two generic search results. Our estimates suggest that distortion of traffic at the first position can be as large as $66\%$ for small arrangement changes. Assuming for the second position the effect is $20\%$ smaller, gives a conservative average effect size of $\beta=0.8\cdot 0.66$. Multiplying the effect size by the fraction of incoming clicks it concerns, we get $0.8\cdot 0.7\cdot\beta\approx 30\%$. This is the fraction of traffic to the site Google can redirect.
@@
Turning this number into a conservative lower bound on performative power, it offers an interpretable measure of power for an investigator to judge whether in this case the algorithmic lever of self-preferencing should be a concern for competition in the down-stream market or not. We can use the same logic to compare search engines, and assess the effectiveness of remedies.
@@
More broadly, we hope our work can serve as a blue print for how performative power can be used to integrate experiments with future digital market investigations, and how tools from computer science, causality, and performative prediction can inform ongoing legal debates related to the power of digital platforms. Our work is situated within a growing scholarship (c.f., Metaxa et al., 2021; Bandy, 2021; Dash et al., 2024) that takes advantage of the accessibility of digital markets for monitoring and regulating digital platforms. Beyond data, we demonstrate how experimental methods can offer an additional tool for power assessments.
@@
From the perspective of computer science our work offers measurements of performativity in the context of online search, contributing quantitative and empirical support to the study of performative prediction. As its name suggests a search engine is performative, it acts as an engine steering consumption through its ranking algorithm, rather than a camera merely picturing candidate results—we borrow this analogy from MacKenzie (2008).